OpenStudy (anonymous):

A student purchases bottled drinks and canned drinks for a party. She purchases a total of 49 drinks for the party at a total cost of $47.55. If bottled drinks each cost $1.15 and canned drinks each cost $.75, How many of each type of drink did she purchase?

OpenStudy (anonymous):

49 drinks = 47.55 Let number of Canned = C, and number of Bottled = B if you wanted to know the cost of a certain amount of bottles and cans you'd use expression such as \[1.15*B = Cost -of- B\] \[.75*C=Cost-of-C\] if we know that the cost of both must equal 47.55, \[1.15B+.75C=47.55\] this is one equation. We also know that the number of bottles and number of cans must equal 49 all together \[B+C=49\]

OpenStudy (anonymous):

use the second to solve the first \[B=49-C\] put this into the first equation and solve for C \[1.15(49-C)+.75C=47.55\]

OpenStudy (anonymous):

i'm going to leave this at this

OpenStudy (anonymous):

if you need more help let me know

OpenStudy (anonymous):

Thank you s much! That helped me a lot